# Quantum Quasinormal Mode Theory for Dissipative Nano-Optics and Magnetodielectric Cavity Quantum Electrodynamics

### v2

@@ VI Conclusion
In conclusion, we have introduced a formalism for the quantization of QNMs in the presence of 3D, spatially inhomogeneous, dissipative (with possible gain), linear magnetodielectric media, enabling a unified and rigorous approach for simulating photonic, plasmonic and magnonic cQED phenomena. Similar to a number of recent works Franke et al. (2019); Medina et al. (2021); Feist, Fernández-Domínguez, and García-Vidal (2020), our approach is rooted in the phenomenological Green’s function quantization scheme of mQED, but relies on the use of PMLs to achieve a set of modes that are rigorously regularized, orthogonalized and complete, rendering them suitable for the representation of the dyadic Green’s function and electromagnetic fields both inside and outside the resonator volume. The introduction of PMLs transforms the radiative losses into non-radiative material dissipation, and via a suitable transformation that reflects all the losses of the resonator, we are able to define creation and annihilation operators that allow the construction of modal Fock states for the joint excitations of field-dressed matter. Our quantized description of the QNMs thereby provides the foundations for a quantum dynamic treatment of magnetodielectric cQED systems, utilizing discrete modes together with rigorously derived quantum Langevin and master equations. By directly addressing the intricacies of modal loss in a fully quantum theory of magnetodielectric cQED, our approach enables a proper assessment of the performance metrics for nanophotonic device protocols in the Quantum 2.0 era, extending from on-demand sources of single photons, plasmons and magnons, dynamic entanglement generation and quantum state transduction, to light-matter strong coupling in hybrid configurations. Moreover, our quantum QNM theory could prove especially valuable in revealing the opportunities that dissipation may offer in manipulating and tailoring the dynamics of quantum systems (contrary to its traditionally perceived adversarial role), such as in the development of quantum technologies based on quantum nanoplasmonic coherent perfect absorption Lai et al. (2024), non-hermitian topological magnonics Yu et al. (2024) and dissipative photonic time crystals Lyubarov et al. (2022); Liu et al. (2023); Sustaeta-Osuna, García-Vidal, and Huidobro (2025). Note that although the modification of mQED in the case of time-dependent material properties is not trivial, important first steps have very recently been made with the proposition of an mQED Lagrangian for a time-dependent Drude model Horsley and Baker (2025), following the quantization of the electromagnetic field in lossless systems Pendry and Horsley (2024).
@@ Acknowledgements.
The authors gratefully acknowledge funding from Research Ireland via Grants No. 18/RP/6236 and 22/QERA/3821.
Data Availability Statement
The data that support the findings of this study are available from the corresponding author upon reasonable request.

### v3

@@ VI Conclusion
In conclusion, we have introduced a formalism for the quantization of QNMs in the presence of 3D, spatially inhomogeneous, dissipative (with possible gain), linear magnetodielectric media, enabling a unified and rigorous approach for simulating photonic, plasmonic and magnonic cQED phenomena. Similar to a number of recent works franke2019quantization; medina2021few; feist2020macroscopic, our approach is rooted in the phenomenological Green’s function quantization scheme of mQED, but relies on the use of PMLs to achieve a set of modes that are rigorously regularized, orthogonalized and complete, rendering them suitable for the representation of the dyadic Green’s function and electromagnetic fields both inside and outside the resonator volume. The introduction of PMLs transforms the radiative losses into non-radiative material dissipation, and via a suitable transformation that reflects all the losses of the resonator, we are able to define creation and annihilation operators that allow the construction of modal Fock states for the joint excitations of field-dressed matter. Our quantized description of the QNMs thereby provides the foundations for a quantum dynamic treatment of magnetodielectric cQED systems, utilizing discrete modes together with rigorously derived quantum Langevin and master equations. By directly addressing the intricacies of modal loss in a fully quantum theory of magnetodielectric cQED, our approach enables a proper assessment of the performance metrics for nanophotonic device protocols in the Quantum 2.0 era, extending from on-demand sources of single photons, plasmons and magnons, dynamic entanglement generation and quantum state transduction, to light-matter strong coupling in hybrid configurations. Moreover, our quantum QNM theory could prove especially valuable in revealing the opportunities that dissipation may offer in manipulating and tailoring the dynamics of quantum systems (contrary to its traditionally perceived adversarial role), such as in the development of quantum technologies based on quantum nanoplasmonic coherent perfect absorption lai2024room, non-hermitian topological magnonics yu2024mag and dissipative photonic time crystals lyubarov2022ptc; liu2023ctc; osuna2025pair. Note that although the modification of mQED in the case of time-dependent material properties is not trivial, important first steps have very recently been made with the proposition of an mQED Lagrangian for a time-dependent Drude model horsley2025macroscopic, following the quantization of the electromagnetic field in lossless systems pendry2024qed.
@@ Acknowledgements.
The authors gratefully acknowledge funding from Research Ireland (formerly Science Foundation Ireland) via Grants No. 18/RP/6236 and 22/QERA/3821.
Conflict of Interest
The authors have no conflicts to disclose.
Data Availability
The data that support the findings of this study are available from the corresponding authors upon reasonable request.